# VideoGen-of-Thought: Step-by-step generating multi-shot video with minimal manual intervention

### v2

@@ 5.1 Experiment Settings
Current video datasets lack sufficient multi-shot narratives with consistent characters across scenes. We therefore constructed a benchmark dataset using VGoT to create ten 30-shot stories (300 shots in total) for evaluation. Each story originates from a single user input $S$, which generates shot outlines $S^{\prime}$ and detailed scripts $\mathcal{P}$ across 30 shots and five domains defined in Eq. 6. For quantitative assessment, we employ four key metrics: Within-Shot Face Consistency ($\Omega_{\text{WS-FC}}$), Cross-Shot Face Consistency ($\Omega_{\text{CS-FC}}$), Within-Shot Style Consistency ($\Omega_{\text{WS-SC}}$), and Cross-Shot Style Consistency ($\Omega_{\text{CS-SC}}$). We additionally report CLIP score [38], PSNR [14], and IS [4]. Our implementation uses multiple GPT-4o [1] for scripting, Kolor [47] as base model for keyframes, and DynamiCrafter [52] for video generation, compared against EasyAnimate [53], CogVideo [27], and VideoCrafter [7, 8] on NVIDIA H100 GPUs.
@@ 5.2 Comparison Evaluation
VGoT achieves 0.8138 $\Omega_{\text{WS-FC}}$ and 0.9717 $\Omega_{\text{WS-SC}}$, outperforming the best baselines (VideoCrafter2’s 0.5569 $\Omega_{\text{WS-FC}}$ and 0.7981 $\Omega_{\text{WS-SC}}$) by 46.1% and 21.7% respectively. For cross-shot consistency, VGoT’s 0.2688 $\Omega_{\text{CS-FC}}$ and 0.4276 $\Omega_{\text{CS-SC}}$ surpass VideoCrafter1’s second-best 0.0686 $\Omega_{\text{CS-FC}}$ and CogVideo’s 0.2069 $\Omega_{\text{CS-SC}}$ by 291.8% and 106.7%. While outperform text-visual alignment via best CLIP score (0.2557), VGoT maintains this performance while requiring 30$\times$ less manual input.
Human evaluations (Table 2) confirm these findings: VGoT receives 66.7% ”Good” ratings for cross-shot consistency versus 27.2% for VideoCrafter2 and 23.2% for CogVideo. In visual quality, 70.0% of evaluators rate VGoT’s outputs as ”Good” compared to 52.4% for CogVideo. This preference is qualitatively validated in Fig 4, which demonstrates VGoT’s superior maintenance of character consistency and visual coherence across extended narratives compared to baseline outputs.
Figure 5: Visual Demonstration of the ablation studies of VGoT
@@ 6 Conclusion
We present VideoGen-of-Thought (VGoT), a structured framework that automates multi-shot video generation with reasonable storylines and visual consistency from one sentence through minimal manual intervention, through three core innovations: dynamic storyline modeling, identity-aware cross-shot propagation, and adjacent latent transition mechanisms. VGoT achieves 20.39% higher within-shot face consistency and 17.36% better style consistency than previous state-of-the-art methods, with over 100% improvements in cross-shot metrics and 10× fewer manual adjustments than alternatives while maintaining director-level narrative flow. Our work redefines automated multi-shot generation, bridging raw visual synthesis with cinematic storytelling through systematic visual stories decomposition.
Supplementary Material
7 Detailed Example of Dynamic Storylines
@@
For example, consider the user input: e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friend Mike and Jane, one bad guy Tom) and their fantasy adventure in an ancient pyramid.. The LLM would first transform this input into a detailed version $S^{\prime}$ consisting of 30 one-sentence scripts like $s_{1}$: “Mike’s Discovery: Mike examines an ancient map with intense focus, revealing the path to the pyramid.” Subsequently, each $s_{i}$ in $S^{\prime}$ is decomposed into prompts $P$, such as $p_{1}$: Character: Mike, holding an ancient map with Jane by his side. Background: A dense jungle filled with mist and towering trees. Relation: Mike studies the map closely, pointing to a pyramid. Camera Pose: Medium shot focusing on Mike and Jane. HDR Description: Soft light filters through the trees, creating dynamic shadows on the map and characters.”. This structured approach ensures narrative coherence across multiple shots, laying the foundation for the generation process.
@@
$p_{2}^{(1)}=\begin{cases}p_{\text{cha}}&\text{Team of scientists with mixed ages}\\ p_{b}&\text{Helicopter interior}\\ p_{r}&\text{Studying maps}\\ p_{\text{cam}}&\text{Close-up}\\ p_{h}&\text{(Missing HDR specification)}\end{cases}$
@@
$p_{2}^{*}=\begin{cases}p_{\text{cha}}&\text{Diverse team in expedition gear, anxious expressions}\\ p_{b}&\text{Helicopter over Amazon rainforest}\\ p_{r}&\text{Team engrossed in equipment}\\ p_{\text{cam}}&\text{Wide shot showing interior/exterior}\\ p_{h}&\text{Sunlight streaming through windows, warm glow}\end{cases}$
@@
8 Additional Results
We validate VideoGen-of-Thought (VGoT) through four narrative archetypes. Type 1: Longitudinal Character Development demonstrates decade-spanning consistency using prompts like “30-shot story of Marco discovering ancient secrets through culinary journeys”, where our framework maintains consistency across aging sequences. Type 2: Multi-Actor Scenes handles complex group dynamics in scenarios like “Abandoned factory transformed into community art center”, preserving relationship continuity between complex stories and multiple characters across shots through identity-aware propagation. Type 3: Non-Human Narratives extends identity preservation to fantastical subjects, as shown in “Mycelium networks and mechanical bees restoring ecosystems”, we explore the creatity of VGoT in marvelous entities. We also evaluate the capability of VGoT to create diverse stories with the same input: “an immigrant’s story of moving to a new country, struggling, and eventually finding success as an entrepreneur.” in Type 4 showcases.
@@ 8 Additional Results
9 User Study

### v3

@@ 5.1 Experiment Settings
Current video datasets lack sufficient multi-shot narratives with consistent characters across scenes. We therefore constructed a benchmark dataset using VGoT to create ten 30-shot stories (300 shots in total) for evaluation. Each story originates from a single user input $S$, which generates shot outlines $S^{\prime}$ and detailed scripts $\mathcal{P}$ across 30 shots and five domains defined in Eq. 6. For quantitative assessment, we employ four key metrics: Within-Shot Face Consistency ($\Omega_{\text{WS-FC}}$), Cross-Shot Face Consistency ($\Omega_{\text{CS-FC}}$), Within-Shot Style Consistency ($\Omega_{\text{WS-SC}}$), and Cross-Shot Style Consistency ($\Omega_{\text{CS-SC}}$) defined in Appendix C. We additionally report CLIP score Radford et al. (2021), PSNR Fardo et al. (2016), and IS Barratt and Sharma (2018). Our implementation uses multiple GPT-4o Achiam et al. (2023) for scripting, Kolor Team (2024) as base model for keyframes, and DynamiCrafter Xing et al. (2025) for video generation, compared against EasyAnimate Xu et al. (2024), CogVideo Hong et al. (2022), and VideoCrafter Chen et al. (2023, 2024) on NVIDIA H100 GPUs.
@@ 5.2 Comparison Evaluation
[…]
Human evaluations (Table 2) confirm these findings: VGoT receives 66.7% "Good" ratings for cross-shot consistency versus 27.2% for VideoCrafter2 and 23.2% for CogVideo. In visual quality, 70.0% of evaluators rate VGoT’s outputs as "Good" compared to 52.4% for CogVideo. This preference is qualitatively validated in Fig 4, which demonstrates VGoT’s superior maintenance of character consistency and visual coherence across extended narratives compared to baseline outputs.
@@ 6 Conclusion
We present VideoGen-of-Thought (VGoT), a structured framework that automates multi-shot video generation with reasonable storylines and visual consistency from one sentence through minimal manual intervention, through three core innovations: dynamic storyline modeling, identity-aware cross-shot propagation, and adjacent latent transition mechanisms. VGoT achieves 20.39% higher within-shot face consistency and 17.36% better style consistency than previous state-of-the-art methods, with 10× fewer manual adjustments than alternatives while maintaining director-level narrative flow. Our work redefines automated multi-shot generation, bridging raw visual synthesis with cinematic storytelling through systematic visual stories decomposition.
Appendix A Limitations, Licenses, and Future Work
Limitations.
VGoT relies on pretrained components without additional finetuning, which bounds performance by the base models’ capabilities. In particular, DynamiCrafter Xing et al. (2025) may limit motion diversity under highly complex camera trajectories, reduce temporal stability in out-of-distribution scenes with rapid appearance changes, and constrain very long-range dependencies beyond adjacent-shot transitions. These constraints are characteristic of training-free pipelines and motivate future model-level improvements.
Licenses Declaration.
We acknowledge and comply with licenses of third-party assets used in our pipeline. DynamiCrafter is distributed under the Apache License 2.0. Kolor is distributed under the Apache-2.0 license. We use these tools within their permitted scopes and cite their sources. Other services used for scripting (e.g., commercial LLM APIs) are accessed under their respective terms of service. Our released code and evaluation scripts will clearly indicate all external dependencies and their licenses.
Future Work.
We plan to: (1) integrate stronger video backbones and optional finetuning to enhance motion diversity and long-range temporal reasoning; (2) extend identity handling to multi-subject IPP with fine-grained attribute disentanglement; (3) broaden cultural and linguistic coverage in prompts and benchmarks; (4) include optional professional and sturctured movie screenplay writing in the script generation process.
Appendix B Detailed Example of Dynamic Storylines
@@
For example, consider the user input: e.g., a set of one-sentence prompts, 30 shots, describe a story about three avatars (two friend Mike and Jane, one bad guy Tom) and their fantasy adventure in an ancient pyramid.. The LLM would first transform this input into a detailed version $S^{\prime}$ consisting of 30 one-sentence scripts like $s_{1}$: “Mike’s Discovery: Mike examines an ancient map with intense focus, revealing the path to the pyramid.” Subsequently, each $s_{i}$ in $S^{\prime}$ is decomposed into prompts $P$, such as $p_{1}$: Character: Mike, holding an ancient map with Jane by his side. Background: A dense jungle filled with mist and towering trees. Relation: Mike studies the map closely, pointing to a pyramid. Camera Pose: Medium shot focusing on Mike and Jane. HDR Description: Soft light filters through the trees, creating dynamic shadows on the map and characters.". This structured approach ensures narrative coherence across multiple shots, laying the foundation for the generation process.
@@
[…]
@@
[…]
@@
Appendix C Multi-Shot Evaluation Protocol
where $F_{j}^{i}$ represents facial features from frame $j$ of shot $i$ extracted using InsightFace Ren et al. (2023); Guo et al. (2021); Gecer et al. (2021); An et al. (2022, 2021); Deng et al. (2020a, b); Guo et al. (2018); Deng et al. (2018, 2019).
Figure 5: Visual Demonstration of the ablation studies of VGoT
Within-Shot Style Consistency (WS-SC) quantifies stylistic coherence through VGG-19 Simonyan and Zisserman (2014) features:
where $S_{j}^{i}$ represents flattened VGG-19 Simonyan and Zisserman (2014) features from frame $j$ of shot $i$.
@@ Appendix C Multi-Shot Evaluation Protocol
Appendix D Additional Results
We validate VideoGen-of-Thought (VGoT) through four narrative archetypes. Type 1: Longitudinal Character Development demonstrates decade-spanning consistency using prompts like “30-shot story of Marco discovering ancient secrets through culinary journeys", where our framework maintains consistency across aging sequences. Type 2: Multi-Actor Scenes handles complex group dynamics in scenarios like “Abandoned factory transformed into community art center", preserving relationship continuity between complex stories and multiple characters across shots through identity-aware propagation. Type 3: Non-Human Narratives extends identity preservation to fantastical subjects, as shown in “Mycelium networks and mechanical bees restoring ecosystems", we explore the creatity of VGoT in marvelous entities. We also evaluate the capability of VGoT to create diverse stories with the same input: “an immigrant’s story of moving to a new country, struggling, and eventually finding success as an entrepreneur." in Type 4 showcases.
Appendix E User Study
@@
Appendix F Ethics Statement
Potential Harms Caused by the Research Process.
Our study uses publicly available pretrained systems for scripting and generation (e.g., GPT-4o for script preparation, Kolor for keyframes, and DynamiCrafter for video synthesis) in accordance with their licenses and terms of service (see Section 5.1). Computation for metric evaluation and visual synthesis was performed on NVIDIA H100 GPUs. The 10-story benchmark used for evaluation is created by the authors specifically for this work; it does not contain personal data or copyrighted media beyond model outputs generated under the respective tools’ usage policies. A small human evaluation with 10 participants was conducted to assess perceived quality and consistency (Figure 8); participants were informed of the study purpose, their privacy was protected, and compensation followed local norms. No sensitive personal information was collected, and we identified no additional risks to participants.
Societal Impact and Potential Harmful Consequences.
VGoT is a training-free pipeline that automates multi-shot video generation from a single sentence. While this can benefit creative workflows and prototyping, risks remain. First, the environmental footprint of generative pipelines is non-negligible; H100-class accelerators consume substantial energy during inference and evaluation. Second, synthetic videos may be misused for disinformation or to fabricate misleading content if deployed irresponsibly. Third, bias can arise from scriptwriting and scene conventions (e.g., English-centric prompts or specific cultural settings), potentially reducing representativeness across regions or languages. Future work should prioritize energy-aware configurations, content provenance indicators, and broader cultural coverage in story prompts and evaluation scenarios.
Impact Mitigation Measures.
We intend to release the VGoT code and evaluation scripts under an open-source license for academic research use, documented to clarify intended use and discourage misuse. We recommend that downstream deployments include visible AI-generation disclosures and optional watermarking, follow model and dataset licenses, and avoid use cases that could harm individuals or communities. We will maintain the released materials and welcome community feedback to improve responsible usage and coverage.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction state the three core contributions and reported gains, which match the methods in Section 4 and our experimental evidence in Sections 5.1 and C.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Appendix A details our limitations tied to base model capacity and discusses future directions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We present a systematic problem–solution framework for multi-shot video generation in Section 4 and empirical evaluation protocol in Appendix C; we do not include formal theorems or proofs that require explicit assumption sets.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Core settings (10 stories × 30 shots), baselines, models, hardware, and metrics are specified in Sections 5.1 and C; we also state intent to release code and evaluation scripts to facilitate exact reproduction (Appendix F).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We intend to release the VGoT code and evaluation scripts for academic research with documentation and licensing guidance (Appendix F).
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify story construction (10 stories, 30 shots each), baselines, models (GPT-4o, Kolor, DynamiCrafter), hardware (H100), and metrics; our pipeline is training-free, so training hyperparameters are not applicable (Section 5.1).
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Despite we conduct quantitative comparison in Table 1 and qualitative comparison in Figure 4, human evaluation in Table 2, and ablation study in Table 3 and Figure 5, we don’t report point estimates without error bars, confidence intervals, or statistical tests.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We disclose compute resources (H100 GPUs) and a training-free inference pipeline; experiments are reproducible under the specified story set and metrics, with compute bounded by inference-time usage (Section 5.1).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We use synthetic prompts and public pretrained models and do not process personal data; we adhere to standard research ethics and anonymization for submission.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our Ethics Statement (Appendix F) discusses potential positive uses and negative societal impacts (energy footprint, misuse risks, and cultural bias) and outlines mitigation measures.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release high-risk models or scraped datasets; our paper does not introduce assets that require special safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Appendix A declares licenses for DynamiCrafter (Apache License 2.0) and Kolor (Apache-2.0) and notes compliance with external services’ terms.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will introduce new pipeline code and evaluation scripts for VGoT and will provide documentation and license notes upon release (See Appendix F).
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Our Ethics Statement summarizes participant privacy protection and compensation norms; Figure 8 illustrates the interface. We provide the instruction in the supplemental material.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our human-subject study involved rating multi-shot video quality and consistency, which is generally considered minimal risk. Participants were informed about the evaluation task; as detailed in Appendix F, we protected privacy and identified no additional risks. We do not provide formal IRB approval details, which can be common for minimal-risk studies depending on institutional policies, but we followed ethical considerations regarding compensation and privacy.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLM usage is a core component of our method and is described in Section 4 (dynamic storyline modeling) and Section 5.1.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.